OpenStudy (anonymous):

Find the extreme values of the function on the interval and where they occur. f(x) = 1x - 2; -3 ≤ x ≤ 4

OpenStudy (phi):

what is your function f(x) = x - 2??

OpenStudy (anonymous):

im not sure. thats all the question says. im supposed to give a minimum and maximum

OpenStudy (phi):

normally people do not write 1x-2 (they just write x-2) so I am suspicious that the function is something else.

OpenStudy (phi):

can you take a screen shot of the question?

OpenStudy (anonymous):

yes hold on

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

@phi wrong question! sorry

OpenStudy (anonymous):

ok heres the queston @phi

OpenStudy (phi):

OK, it is probably a goofy equation generated automatically by a computer. to find the extreme values of f(x) = 1x - 2; -3 ≤ x ≤ 4 (or any function) you do three things: 1. evaluate f(x) at the lower limit 2. evaluate f(x) at the upper limit 3. take the derivative, and solve for when equal to 0 for step 3: f'(x) = 1 which is never equal to 0. so the extreme values will only be from steps 1 and 2.

OpenStudy (phi):

of course, you do not need calculus for this problem. It is a straight line going up from left to right... so the min value will be at the left side of the interval (at x=-3) and the max value will be when x=4

OpenStudy (anonymous):

ok so of those answers its b?

OpenStudy (anonymous):

@phi i got b "Maximum is 6 at x = 4; minimum value is - 1 at x = -3"

OpenStudy (phi):

f(x) = 1x - 2 at x=-3 we get -3-2 = -5 at x= +4 we get 4-2 = 2 you must have done something different from the steps I gave in the other post.

OpenStudy (anonymous):

ok i was redoing my work and found what i did wrogn. thanks